# Evaluating Cooling Center Coverage Using Persistent Homology of a Filtered Witness Complex

### v1

@@ 3.1 Our Construction of the Witness Complex
Details on the data we use can be found in Appendix A.1 and A.2. Our code for constructing the witness complex can be found on GitHub [^1] https://github.com/erinponeil/Cooling_Center_Converage_TDA.git. We use the GUDHI library [30] for computing persistent homology.
@@
3.2 Our Heat Vulnerability Index Score
@@
To test our method against more standard approaches in this field, we compute a heat vulnerability index (HVI). This is based on characteristics of the demographics and environment that have been shown to exacerbate the likelihood that a person will suffer negative health consequences from an extreme heat event. Due to data availability, this score is defined on census tracts, not census blocks. The characteristics of a census tract $T$ that we use are the typical afternoon temperature $F(T)$ in Fahrenheit, the percentage of the area not covered by tree canopy $C(T)$, and the number of younger and older residents, $Y(T)$ and $S(T)$ respectively [^2] Note that the National Integrated Heat Health Information System defines “young” to mean under 5 years old and “older” to mean over 65 years old.. The HVI score $V(T)$ of census tract is
(2) $V(T)=\frac{F(T)-\mu_{F}}{\sigma_{F}}+\frac{C(T)-\mu_{C}}{\sigma_{C}}+\frac{Y(T)-\mu_{Y}}{\sigma_{Y}}+\frac{S(T)-\mu_{S}}{\sigma_{S}}\,,$
where $\mu_{F}$ is the mean and $\sigma_{F}$ is the standard deviation of temperatures across all census tracts within a given city (other variables are defined analogously). This unweighted $z$-score approach is modeled after other HVI computations [5, 6, 21]. We standardize the variables before creating a composite score to account for their differing units of measurement.
Various factors influence an individual’s risk of heat-related mortality. Our motivation for incorporating these four variables into our measure of vulnerability is detailed as follows. Firstly, one can imagine that a warmer census block with few trees may not fare well during an extreme heat event. Secondly, being over the age of 65 puts an individual more at risk of heat related mortality due to social isolation and a lack of mobility to get assistance when faced with emergencies [16]. Lastly, because children thermoregulate less effectively than adults, they are more likely to respond poorly during extreme heat events [16]. The census block is interpreted to be at greater risk when its HVI score is higher. See Figure 3 for the maps of HVI scores across the four regions of interest. Cutter et al. explains that an essential component of building a HVI score involves testing for multicollinearity among demographic variables [8]. We verified that the four variables used in this study are independent using variance inflation factor (VIF). More details on the demographic data used to create the HVI score can be found in Appendix A.3. Refer to Table 2 for the results of the VIF analysis.
Figure 3: Heat vulnerability index (HVI) score maps for the four regions of interest. Hatching highlights census tracts that had one or more missing data value(s) for typical afternoon temperature, area not covered in by tree canopy, the number of younger residents, or the number of older residents.
4 Results
We test our methodology on four locations in the United States that offer unique climates and demographics. Those locations include central Boston, MA; central Austin, TX; Portland, OR; and Miami, FL. These cities differ in various ways, for example Miami and Boston are on the water, while Austin is landlocked. Additional differences include humidity, which can increase how warm a temperature feels, and the populations of vulnerable demographics such as older adults. It is important to note that all the selected cities are vulnerable to the urban heat island effect [26]. In the context of our data, we are able to deduce which census blocks (landmarks) in our simplicial complex remain disconnected from other census blocks for a longer duration. This can be interpreted as more “vulnerable” as they are further away from cooling centers compared to other census blocks. We compute the PH of the filtered witness complex for each of the four locations of interest. Their persistence diagrams (PDs) are shown in Figure 4. Notably, the PD for Boston shows significantly more 1D homology class points than any other location.
@@ 4 Results
Figure 4: Persistence diagram for our four geographic locations. Note that the axes are not on the same scale.
Figure 5 maps the death simplices of the 0D homology group for our four geographic locations. Those with larger death values are interpreted as connecting one region to one with higher risk of heat-related mortality. For instance, the southern tip of Austin (census tract 24.32) has an especially large death time. Other high death values can be observed in the south-east area of central Boston, on the northern (coastal) tip and south-west (coastal) edge of Miami, and the north-west and south-east tip of Portland.
@@
Figure 5: Map of death values for the 0-dimensional homology classes. Largest death values are represented in yellow. Maps are not to scale.
We identify the components (landmarks) that correspond to the persistence points with the five largest death times for each city. These centroids are marked with stars in Figure 6 colored based on the death time. Our topological approach identifies locations that are mostly on the edge of the city despite us adding cooling centers outside the city limits. Figure 6 also shows the census tracts with the five highest HVI scores in each city. Considering these vulnerable areas, we can suggest optimal locations for the development of future cooling centers to better serve those at risk. It can be observed that these suggested locations are frequently in areas that currently lack cooling centers. The suggestion of additional cooling centers on the eastern edge of central Austin (census tract 22.07), for example, may prove beneficial considering it has nearly 10,000 residents of whom 1,100 are of a vulnerable age to heat related mortality. Similarly, additional cooling centers in the south-east tip of Portland could cover those residents living in census tracts of high HVI scores, as seen in Figure 3.
Figure 6: These maps include the distribution of cooling centers, the locations considered to have the highest vulnerability according to our TDA approach due their large death times (stars), and the top five census tracts with the largest HVI scores (grey regions). Our approach provides a complimentary set of data points to those uncovered by the HVI map.
Figure 7 maps the death simplices the 1D homology group for our four geographic locations. Death simplices with higher death value can be interpreted as a “hole” in coverage. For instance, high death values can be observed at the south-west edge of Boston and can be viewed as a high risk neighborhood during an extreme heat event.
@@
Figure 7: Map of death values for the 1-dimensional homology classes. Largest death values are represented in yellow. Maps are not to scale.
Figure 8 summarizes the 0D and 1D death values for the four locations of interest. While the differences in death times are not statistically significant, Austin has many more outliers with high death time than any other city, possibly indicating greater risk.
Figure 8: This box plot illustrates the death values of the 0D and 1D homology classes for each of the four geographic locations of interest.
5 Conclusion and Discussion
@@
In this paper, we use persistent homology to determine which locations are most in need of cooling centers to reduce mortality risk. Our methodology presents some advantages over traditional approaches for analyzing cooling center coverage. Our TDA approach only requires the current distribution of cooling centers and the centroids of subdivisions (e.g. census tracts or census blocks) within a given city. Demographic data, which can be difficult to collect, is not needed. Although TDA is not a replacement to standard techniques such as an HVI map, it can identify a complementary set of high risk locations which could further inform policymakers of the ideal placement of additional cooling centers. Additionally, our approach does not treat each block as an isolated “island.” Instead, it provides a holistic evaluation of the spatial distribution of cooling centers and considers neighboring cities, adjacent blocks, and the proximity of those blocks to cooling centers.
@@ 5.1 Limitations
Similarly, we are limited by the completeness of the demographic data. For instance, we restricted our attention to central Austin, TX and central Boston, MA because the U.S. Urban Heat Island Mapping Campaign only included data in that region. In addition, ideally we would have demographic data at the census block level rather than the census tract level. Moreover, some census tract scores may be lower or higher in reality due to missing data. For instance, the data for census tract 9812.02 in Boston, MA showed no residents were less than 5 or over 65, which is attributed to missing data.
Lastly, it is important to address the topographical features that may limit our analysis. For one, census blocks can be different sizes within a given city. This is a problem, for example, because a centroid in a large census block may appear to be isolated spatially in our analysis and, therefore, not connect to other centroids until later in the filtration. This census block could be labelled as not being well covered, when in fact it is. Ideally, a census block’s centroid would be a good geographical approximation for any neighborhood within the census. In addition, our analysis does not account for bodies of water. For example, some unpopulated islands off the coast of Boston are included within census block boundaries without regards to the ocean.
5.2 Future Work
Future work includes adapting additional techniques used by Hickok et al. [19], such as measuring distance in travel time and incorporating several modes of transportation. Google maps has a fee for each distance calculated, so using travel time or different modes of transportation was infeasible in this case. The geodesic distance has a lower computational and monetary cost.
In the future, we would like to conduct a larger scale investigation about additional cities. We chose four different cities to test our method, however there are many other cities included in the US Urban Heat Island Mapping Campaign. Our methodology could be used in many other applications as well; for example, studying the coverage of resources such as food banks or healthy food options. On a smaller scale, one could consider the locations of automated external defibrillators (AEDs) or fire extinguishers within a large building, such as a school. A resource that will be important over the coming decades is electric vehicle charging stations. For this application, a larger statewide or nationwide scale would be useful to determine areas where the addition of a charging station would facilitate the use of electric vehicles in the area.

### v2

@@ 3.1.1. Interpretation and Visualization
Details on the data we use can be found in Appendix A.1 and A.2. Our code for constructing the witness complex can be found on GitHub. [^6] https://github.com/erinponeil/Cooling_Center_Converage_TDA.git We use the GUDHI library [35] for computing persistent homology.
3.1.2. Distance Computation
We consider the fact that each city is in a different region of the country. This is an important consideration because distances, like a Euclidean distances, can quickly become inaccurate due to the failure to account for the curvature of the Earth. In other words, a projected map of the US may distort distances in some cities more than others. To avoid the distortion of projected maps, we compute distances between latitude and longitude points using a straight-line measurement along Earth’s surface which is often referred to as the geodesic distance. We believe this choice to be justifiable because some studies have shown straight-line measurements are an acceptable proxy to travel distances in urban areas [3, 17, 26]. Furthermore, in non-emergency situations, Boscoe et al. concluded that the difference between straight-line and driving estimates were inconsequential and, because of this, straight-line estimates are often used [17, 32]. This measure of distance is then converted into units of kilometers for easier interpretation.
@@ 3.1.2. Distance Computation
Figure 3. Heat vulnerability index (HVI) score maps for the four regions of interest.
3.2. Heat Vulnerability Index (HVI)
In order to compare our topological approach against standard methods for measuring heat vulnerability, we construct a Heat Vulnerability Index (HVI) based on the existing literature. Multiple methods exist for calculating HVIs in urban areas, including principal components analysis (PCA), ordered weighted averaging (OWA), hierarchical agglomerative clustering, theoretical component (TC), and unweighted $z$-scores, among others [2, 29]. For simplicity, we opt to use an unweighted $z$-score that is modeled after several other HVIs [7, 8, 25]. In this study, the HVI map serves as a contextual overlay and comparison for our TDA analysis, providing insight into the demographics of the city under investigation. [^7] One can construct their own contextual HVI maps using alternative variables, methods, and weights using the code in https://github.com/erinponeil/Cooling_Center_Converage_TDA.git.
@@ 3.2. Heat Vulnerability Index (HVI)
The HVI is based on characteristics of the demographics and environment that have been shown to exacerbate the likelihood that a person will suffer negative health consequences from an extreme heat event. Since the demographic data is not available at the census block level, we define this score on census tracts. The characteristics of a census tract $T$ that we use are the typical afternoon temperature $F(T)$ in Fahrenheit, [^8] Note that afternoon temperature data was collected at 3pm in the respective city using a combination of Sentinel-2 satellite data and sensors mounted on the bikes and cars of volunteers [31] the percentage of the area not covered by tree canopy $C(T)$, and the number of younger and older residents, $Y(T)$ and $S(T)$ respectively. [^9] Note that the National Integrated Heat Health Information System defines “younger” to mean under 5 years old and “older” to mean over 65 years old [31]. Because these variables have different units of measurement, we standardize them before creating a composite score. The HVI $HVI(T)$ of census tract $T$ is
(2) $HVI(T)=\frac{F(T)-\mu_{F}}{\sigma_{F}}+\frac{C(T)-\mu_{C}}{\sigma_{C}}+\frac{Y(T)-\mu_{Y}}{\sigma_{Y}}+\frac{S(T)-\mu_{S}}{\sigma_{S}}\,$,
where $\mu_{F}$ is the mean and $\sigma_{F}$ is the standard deviation of temperatures (other variables are defined analogously) across all census tracts in all cities studied in [31]. Because the standardization is done across all cities, we can compare the HVIs across cities.
Various factors influence an individual’s risk of heat-related mortality. We select four variables that span demographic, geographic, and topographic features of the area. See Appendix A.3 for additional discussion of the HVI variables. In particular, Figures 1215 show maps of each of the four variables individually before they are combined to make the HVI shown in Figure 3.
3.2.1. Evaluating Multicollinearity
Cutter et al. explains that an essential component of building a HVI involves testing for multicollinearity among demographic variables [10]. We examine the four variables used in this study and test for independence using variance inflation factor (VIF) [22]. The full VIF results can be found in the Appendix in Table 4, however we briefly discuss them here. The VIF analysis for Austin, Miami, and Boston indicate negligible multicollinearity for all four variables. The results for Portland indicate the afternoon temperature and tree canopy have non-negligible multicollinearity and thus are likely not independent of each other. Because of this, the HVI results for Portland should be considered with caution. We further analyze the independence and multicollinearity using pairwise correlations in Appendix A.3 and Figures 1619.
@@ 3.2.1. Evaluating Multicollinearity
Figure 4. Persistence diagram for our four geographic locations. The birth and death times are in units of kilometers. Note that the axes are not on the same scale.
3.2.2. Limitations
While the purpose of the HVI is to serve as a comparison to our topological approach, we will briefly discuss its limitations. We limit our HVI to only include four variables for the sake of simplicity, however, the U.S. Urban Heat Island Mapping Campaign dataset contains sever other variables that could be incorporated. This includes features such as average morning and evening temperatures, the number of individuals living in poverty, the median income, and the percentage of the population that is a minority [^10] Defined by the U.S. Urban Heat Island Mapping Campaign to mean non-white.. Because we chose two demographic variables, it is possible the HVI results are biased towards places with a higher population. For example, places that are cooler but have a higher population of individuals <5<5 years old or >65>65 years old may have a higher HVI than a place that is warmer but has a lower population of vulnerable individuals. Beyond which variables are included, there are modifications that could be made to the approach. While the unweighted $z$-score approach is standard in the existing literature, it weights each variable equally and it increases linearly with the increase of any one variable. Because measuring heat vulnerability is multifaceted, it is unclear if these are valid assumptions. [^11] One can explore the HVI and incorporate different weightings and combinations of variables using code in our Github repository.
Figure 5. Box plots for each of the four geographic locations of interest of (a) death times of the 0D, (b) death times of 1D homology classes, and (c) the lifetimes of the 1D homology classes. A tables of the corresponding means and standard deviations can be found in Table 1.
Table 1. Mean and standard deviation (SD) of: dimension 0 death times, dimension 1 death times, and dimension 1 lifetimes.
0D Death Time 1D Death Time 1D Lifetime
City Mean SD Mean SD Mean SD
Austin 1.939 1.428 2.243 0.718 0.060 0.041
Boston 0.745 0.454 1.08 0.554 0.144 0.172
Miami 1.15 0.518 1.345 0.609 0.167 0.160
Portland 1.418 0.796 2.425 0.631 0.258 0.343
4. Results
In this section we will analyze the results of the HVI and our topological approach in Sections 4.1 and 4.2 respectively, followed by a comparison of the two approaches in Section 4.3.
We test our methodology on four locations in the United States that offer unique climates and demographics. Those locations include central Boston, MA; central Austin, TX; Portland, OR; and Miami, FL. These cities differ in various ways, for example Miami and Boston are on the water, while Austin is landlocked. Additional differences include humidity, which can increase how warm a temperature feels, and the populations of vulnerable demographics such as older adults. It is important to note that all the selected cities are vulnerable to the urban heat island effect [31].
4.1. HVI
Based on the HVI maps in Figure 3, we can identify the regions of higher HVI which indicate regions more vulnerable to heat related mortality. In particular, we can see individual census tracts in central Austin, northern Boston, western Miami, and eastern Portland with particularly high HVIs. We can see based on the box plots of the HVI values in Figure 6 that all cities have similar medians, although the HVI values in Miami are slightly higher. A table of statistics of the HVI values can be found in Appendix A.3 in Table 5. Additionally, Figures 1215 show heatmaps of each of the individual variables used in the HVI computation. These can be used to identify regions where one particular indictor of vulnerability may be high even if the composite HVI is not.
Figure 6. Boxpot of HVI across all census tracts within a given city.
4.2. Topological Approach
We compute the PH of the filtered witness complex for each of the four locations of interest. Their persistence diagrams (PDs) are shown in Figure 4. Notably, the PD for Boston shows significantly more 1D homology class points than any other location. Figure 5 plots the distributions of 0D and 1D death values for the four locations of interest. The average death times (which are listed in Table 1) are fairly similar across cities in both dimension 0 and 1; the average death time in dimension 0 is between 0.5 and 2 while the average in dimension 1 is between 1 and 2.5. It is worth noting that Austin has many more outliers with high 0-dimensional death time than any other city, possibly indicating more regions with greater risk.
For dimension 0, all features have a birth time of 0 so the death time is also the lifetime. For dimension 1, however, the birth time is nonzero so we also compute the lifetime of the persistence points. The birth time alone is not informative in the context of this application, however if the lifetime is small, it indicates that the hole fills in soon after it is formed so it may not be a significant hole in coverage. Even though in dimension 1 Austin has the second highest average death time, the average lifetime is the smallest across all four cities, likely indicating the regions indicated by the 1-dimensional death simplices are less vulnerable in Austin than in other cities. On the other hand, Boston has the largest death times and lifetimes of 1-dimensional features across all four cities, likely indicating there are more significant holes in coverage in Boston than in Austin.
@@ 4.2. Topological Approach
Figure 7. Death simplices for the 0-dimensional homology classes. Maps are not to scale and the scale of death values are different across the cities.
4.2.1. Death Simplices
As discussed in Section 3.1.1, we use the death simplices to visualize the regions of vulnerability. Figures 7 and 8 map the death simplices of the 0- and 1-dimensional homology features respectively.
In dimension 0, death simplices indicate a stretch of space along which there is a lack of cooling center coverage. Those with larger death values are interpreted as connecting one region to one with higher risk of heat-related mortality. For instance, in Austin, the edge connecting the southernmost census block to the center of the city has an especially large death time. Similarly, there are several death simplices with a high death time on the eastern side of the city. This indicates that the southern and eastern portions of the cities lack coverage and are therefore those who reside in those areas are vulnerable to extreme heat. Other death simplices with high death times can be observed in the south-east area of central Boston, on the northern tip and south-eastern edge of Miami, and the north-west and south-east tip of Portland.
In dimension 1, the death simplices indicate the epicenter of the hole in coverage of cooling centers. In Figure 8, we see one or more death simplices with a high death time in the southern regions of Boston and Portland as well as in central Miami. These regions are the most vulnerable to heat related mortality. Boston also has numerous death simplices with greater-than-average death values. This indicates Boston has more significant holes in coverage than the other three cities.
@@ 4.2.1. Death Simplices
Figure 8. Death simplices for the 1-dimensional homology classes. Largest death values are represented in yellow. Maps are not to scale.
Figure 9. Comparison of the two measures of vulnerability. The vertical axis plots the maximum death time of components containing any census block within a given census tract. The horizontal axis plots the corresponding HVI for the census tract. Table 2 shows the correlation and corresponding $p$-value for each city individually as well as all four combined.
Figure 10. In each map, census tracts are colored based on their HVI value relative to the average HVI within the city. Therefore pink and green tracts have an HVI above and below average, respectively. The blue lines and triangles are the death simplices for dimension 0 and 1 with death times in the 90% percentile of death times within that city.
4.3. Comparing Approaches
The HVI and topological approaches each identify regions of high vulnerability. Here we compare the regions identified by each approach.
For a simple initial comparison, for each census tract we compute the maximum death time of any census block in that tract and plot it against the HVI of that tract in Figure 9. It is hard to see any patterns visually, so we compute the correlations of each city individually as well as all four together. These correlations are reported in Table 2. Only Boston has a statistically significant correlation with a correlation coefficient of 0.132. However, based on the large $p$-values across the other cities, we cannot draw any significant conclusions.
Next we want to look at specific regions. To do so, we plot the death simplices with a death time in the 90-th percentile and the HVI data on the same plot in Figure 10. It is important to note the colorbar in Figure 10 does not represent the HVI. Instead, for a given census tract $T$, it plots $HVI(T)-\text{mean}\{HVI(t):t\in\mathcal{T}\}$ where $\mathcal{T}$ is the set of all census tracts within a given city. Therefore the colorbars should be interpreted as a measure of how vulnerable a tract is relative to the average. Census tracts colored pink (which corresponds to positive values) indicate census tracts with an HVI above the mean. Therefore, any intersection between a death simplex and a pink (above average HVI) census tract indicate regions of agreement between the two measures of vulnerability. We will focus on analyzing this plot to compare when the two methods align and analyze when they differ.
In Austin, we see significant overlap of the death simplices for dimension 0 with the regions of above-average HVI. In particular, on the northeastern portion of Austin, there are two census tracts with well above average HVI that also intersect with death simplices. This indicates agreement between the approaches meaning the regions with high HVI also lack cooling center coverage. Because both methodologies agree in these locations, we can interpret these locations a being more severely vulnerable than the locations identified by only one approach.
In Boston, the northeastern and southeastern corners of the city contains numerous death simplices and a collection of census tracts with above average HVI, indicating regions of high vulnerability. Similarly, in eastern Portland there are several regions that are identified as vulnerable by both approaches, indicating another area of vulnerability.
We can also use this plot to identify regions of disagreement where one approach may identify the region as vulnerable while the other does not. These regions would indicate a moderate level of vulnerability For example, regions with above average HVI but no death simplices could be due to a significant vulnerable population but in an area that is well covered by cooling centers. Regions with a below average HVI but several death simplices could indicate areas that have lower temperatures and more tree cover but do not have many nearby cooling centers. An HVI well below average with death simplices might indicate a low-population area where cooling centers are therefore not as necessary. We see examples of all of these in Figure 10.
For example, in Figure 7, we see one death simplex in Boston with a particularly high death time on the eastern side of the city. Based on this we would interpret this region has being highly vulnerable, however this particular census block is located on an island in Boston Harbor called Moon Island. It was previously a sewage treatment plant but now is used as a training facility for the Boston fire and police departments [37]. While it is true this area has poor access to cooling centers, it has a below average HVI, indicating it is not an area of concern. We can investigate further and see in Figure 13 that the census tract has a very low population of individuals of age <5<5 or >65>65. Similarly, in northwest Portland, there is a death simplex with high death time but the corresponding area has a low HVI due to a low number of vulnerable individuals.
Conversely, in south Boston, there are a collection of census tracts with lower-than-average HVI but where there are several death simplices in the area. Consulting Figure 13, we see the area has low average temperatures and a high percentage of tree cover, indicating lower vulnerability, but there is a fairly high number of individuals of age <5<5 or >65>65. So while the HVI would not indicate this as an area of high priority, the lack of cooling centers makes it more concerning.
5. Conclusion and Discussion
In this paper, we use persistent homology to supplement existing approaches for identifying areas with high risk of heat related mortality. In particular, we develop an approach for identifying locations in need of cooling centers to reduce mortality risk. When analyze together with a HVI, a standard approach, we see a more comprehensive understanding of heat-related vulnerability.
@@ 5. Conclusion and Discussion
Our TDA approach only requires the current distribution of cooling centers and a collection of landmarks (e.g. the centroids of census tracts or census blocks) within a given city. Demographic data, which can be difficult to collect, is not needed. However, when used independent of the HVI, we see the topological approach may over or underestimate the vulnerability of a given area. Thus TDA is not a replacement to standard techniques such as an HVI map, but it can identify additional high-risk locations and aid in prioritizing existing high-risk areas. Our topological approach captures the spatial distribution of cooling centers and considers neighboring cities, adjacent blocks, and the proximity of those blocks to cooling centers. In conjunction with the HVI, it provides a more holistic evaluation of the vulnerability than can be identified using either approach individually.
@@ 5.1. Limitations
Similarly, we are limited by the completeness of the demographic data. For instance, we restricted our attention to central Austin, TX and central Boston, MA because the U.S. Urban Heat Island Mapping Campaign did not have comprehensive data for the full city. In addition, ideally we would have demographic data at the census block level rather than the census tract level.
Further, as discussed in , we use the geodesic distance between census blocks and cooling centers. While some work suggests this is an appropriate proxy for travel distance, it cannot account for variables that influence travel such as traffic and road closures. Ideally we would like to use travel time like the approach in [23], however google maps has a fee for each distance calculated, so using travel time or different modes of transportation was infeasible in this case.
Lastly, it is important to address the topographical features that may limit our analysis. For one, census blocks can be different sizes within a given city. This is a problem, for example, because a centroid in a large census block may appear to be isolated spatially in our analysis and, therefore, not connect to other centroids until later in the filtration. This census block could be labeled as not being well covered, when in fact it is. Ideally, a census block’s centroid would be a good geographical approximation for any neighborhood within the block. A future improvement could include sampling multiple points within a census block rather than only using the centroid. In addition, our analysis does not account for bodies of water. For example, some unpopulated islands off the coast of Boston are included within census block boundaries without regards to the ocean.
Table 2. Correlation between death time and HVI, as shown in Figure 9. The last row labeled “All” indicates the correlation if we combined the points for all four cities. Note that Boston is the only city with a $p$-value <0.05<0.05 with a correlation of 0.132, indicating a slightly positive correlation.
City Correlation Coeff. p-value
Austin 0.096 0.446
Boston 0.132 0.018
Miami -0.140 0.171
Portland -0.075 0.372
All 0.048 0.234
5.2. Future Work
In the future, we would like to explore other spatially aware approaches for detecting heat vulnerability from the literature as additional points of comparison. We’d also like to conduct a larger scale investigation about additional cities. We chose four different cities to test our method, however there are many other cities included in the US Urban Heat Island Mapping Campaign.
Our methodology could be used in many other applications as well; for example, studying the coverage of resources such as food banks or healthy food options. On a smaller scale, one could consider the locations of automated external defibrillators (AEDs) or fire extinguishers within a large building, such as a school. A resource that will be important over the coming decades is electric vehicle charging stations. For this application, a larger statewide or nationwide scale would be useful to determine areas where the addition of a charging station would facilitate the use of electric vehicles in the area.